You have been called to testify as an expert witness in a trial involving a head-on collision. Car A weighs 1515 lb and was traveling eastward. Car B weighs 1125 lb and was traveling westward at 41.0 mph. The cars locked bumpers and slid eastward with their wheels locked for 21.5 ft before stopping. You have measured the coefficient of kinetic friction between the tires and the pavement to be 0.750. How fast (in miles per hour) was car A traveling just before the collision? (This problem uses English units because they would be used in a U.S. legal proceeding.)
Answer #1

d = 21. 5 ft =

using work energy theorem to find the velocity of two car systems

u ( m1 + m2) g d = 0.5* (m1 + m2) v^2

0.75* 9.8* 4.07*10^-3 = 0.5* v^2

v = 0.2446 mph

using conservation of linear momentum

m1 v1 + m2 v2 = (m1 + m2) v

1515* :v1 - 1125 * 41 = (1125 + 1515)* 0.2445

v1 = 30.872 mph

=========

Comment in case any doubt, will reply for sure Goodluck